A sample of solid aluminum is heated with an electrical coil. If 250. Joules of energy are added to a 14.7 gram sample initially at 20.5°C, what is the final temperature of the aluminum?
1
Expert's answer
2020-11-10T14:05:27-0500

Q=mc∆T. ∆T=Q/mc = 250/14.7 x 0.887 = 19,1733965.

∆T=T2-T1.

T2= ∆T+T1 = 19,1733965 + 20.5 = 39,6733965 °C.